ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 202, 2014

Apter, A. W. and M. Gitik On tall cardinals and some related generalizations

343

Bachoc, C., E. DeCorte, F. M. Oliveira Filho and F. Vallentin Spectral bounds for the independence ratio and the chromatic number of an operator

227

Barré, S. and M. Pichot Removing chambers in Bruhat–Tits buildings

117

Bergner, J. E. and P. Hackney Group actions on Segal operads  423

Boltje, R. and B. Külshammer Central idempotents of the bifree and left-free double Burnside ring

161

DeCorte, E. See C. Bachoc Demarche, C. and D. Wei Hasse principle and weak approximation

for multinorm equations

275

García-Pacheco, F. J. The universal renorming  405

Gille, S. and N. Semenov On the Brauer group of the product of a torus and a semisimple algebraic group

375

Gitik, M. See A. W. Apter Globk, W. On the geometry of flat pseudo-Riemannian homogeneous

spaces

255

González-Sánchez, J. and A. Zugadi-Reizabal A characterization of powerful p-groups

321

Grahl, J. and S. Nevo Exceptional functions wandering on the sphere and normal families

21

Hackney, P. See J. E. Bergner

Hua, B. and J. Jost L^{q} Harmonic functions on graphs  475

Huang, Y. and S. Wu On periodic Teichmüller disks of modular transformations

491

Jost, J. See B. Hua Külshammer, B. See R. Boltje Kaplan, I. and S. Shelah A dependent theory with few indis-

cernibles

59

Kerr, D. Bernoulli actions of sofic groups have completely positive entropy

461

Kojman, M., D. Milovich and S. Spadaro Noetherian type in topological products 195 Liu, X. Y. and Z. H. Liu Relaxation control for a class of evolution

hemivariational inequalities

35

Liu, Z. H. See X. Y. Liu Milovich, D. See M. Kojman Nevo, S. See J. Grahl Oberlin, D. M. Exceptional sets of projections, unions of k-planes

and associated transforms

331

Peng, C. and X. Xu Minimal two-spheres with constant curvature in the Complex Grassmannians

1

Pichot, M. See S. Barré Semenov, N. See S. Gille Shelah, S. See I. Kaplan Spadaro, S. See M. Kojman Tong-Viet, H. P. finite groups whose irreducible Brauer characters

have prime power degrees

295

Vallentin, F. See C. Bachoc Wei, D. See C. Demarche Wu, S. See Y. Huang Xu, X. See C. Peng Zamora-Aviles, B. Gaps in the poset of projections in the Calkin

algebra

105

Zugadi-Reizabal, A. See J. González-Sánchez de Oliveira Filho, F. M. See C. Bachoc